Please help. Solve for the length and width.
The length of a rectangle is twice the width. The area of the rectangle is 162yd^2. Find L and W.
Thank you

Answer by checkley75(3666)

L=2W
LW=162
2W*W=162
2W^2=162
W^2=162/2
W^2=81
W=SQRT81
W=9 ANS.
L=2*9=18 ANS.
pROOF:
9*18=162
162=162